OpenStudy (anonymous):

whts (14-3C)+7C=94. PLZ

OpenStudy (anonymous):

14-3c+7c=94 Let's start by combining the Cs: -3c + 7c = 4c, so now we have: 14+4c=94 Now we're going to work on getting c alone. To do that we're going to subtract the 14 from that side (and the other side as well so it's still equal): 14-14+4c=94-14 4c=80 We have 4timesC equals 80 and we still need to get the c alone, so to undo the multiplication (4c) we're going to divide the 4 off and, again, we have to do the same to both sides: (4c)/4=80/4 c=20 (because 80 divided by 4 is 20.)

OpenStudy (anonymous):

wait i dont get the first step

OpenStudy (anonymous):

{14-3c}+7c=94=> 14+4c=94=> 4c=94-14=> c=80/4=> c=20

OpenStudy (anonymous):

oh thannnkkkk uuu uuuu

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

now sorry how did u do the 2nd step i get everythin after tht just where did u get the 4c, confused im sorry i might be a bother

OpenStudy (anonymous):

(14 - 3c) + 7c = 94 Add -3c and 7c 14 + 4c = 94 Subtract 14 from both sides 4c = 80 Divide both sides by 4 and the answer is c = 20

OpenStudy (anonymous):

oh finaly thankkkkkkkkkkkkkkkkkk uuuu u made my life easy. Nice takin to \ya.

OpenStudy (anonymous):

np and nice talking to you too